# CLAPP: Contrastive Language-Audio Pre-training in Passive underwater vessel classification

### 04 Jan 2024 v1

@@ 5.1 Comparison with State-of-the-Art Methods
In this subsection, we analyze the effectiveness of each component in the CLAPP framework in zero-shot task. Here, we adopt the CLIP-ViT-B/16 model fine-tuned with SDM loss as the Baseline to facilitate the ablation study.
Ablations on auxiliary information. To fully demonstrate the impact of different components in CLAPP, we conduct a comprehensive empirical analysis on two public underwater audio datasets, deepship and shipsear. The accuracy (%) are reported in Tab. 3.
The text modality of our proposed pretraining dataset, Oceanship, encompasses three distinct types of auxiliary information: location information, ship’s speed over ground, and timestamp. By incorporating these auxiliary information into the text modality, our comprehensive analysis from the table demonstrates that the augmentation of auxiliary information positively correlates with improved performance in the zero-shot classification task, specifically on the Deepship and Shipsear datasets.

### 15 Jan 2024 v2

@@ 5.1 Comparison with State-of-the-Art Methods
In this subsection, we further analyze the effectiveness of each component in the CLAPP framework in zero-shot task. Here, we adopt the CLIP-ViT-B/16 model fine-tuned with SDM loss as the Baseline to facilitate the ablation study. More detailed information can be listed as follows:
Ablations on auxiliary information. To fully demonstrate the impact of different components in CLAPP, we conduct a comprehensive empirical analysis on two public underwater audio datasets, named deepship and shipsear respectively. More detailed accuracy (%) performance are reported in Tab. 3.
In general, the text modality of our proposed pretraining dataset, Oceanship, encompasses three distinct types of auxiliary information: location information, ship’s speed over ground, and timestamp. According to the experimental results in Tab. 3, by incorporating these auxiliary information into the text modality, our comprehensive analysis from the table demonstrates that the augmentation of auxiliary information positively correlates with improved performance in the zero-shot classification task, specifically on the Deepship and Shipsear datasets.